# Markov spin models for image generation : explicit large deviations with respect to the number of pixels

### v1

@@ IV.3 Large deviations of the backward-generated distribution of the intensive overlap $q$ for large $N$
For large $N$, using the Stirling approximation of Eq. 35, one obtain that the large deviations of the intensive overlap $q=\frac{Q}{N}$ in the backward-generated distribution of Eq. 109
@@
$i_{t}^{Gene}(q)$ $=\frac{1+q}{2}\ln\left(\frac{1+q}{\frac{(1+\lambda^{t})(1-\lambda^{2T-t})}{(1-\lambda^{2T})}}\right)+\frac{1-q}{2}\ln\left(\frac{1-q}{\frac{(1-\lambda^{t})(1+\lambda^{2T-t})}{(1-\lambda^{2T})}}\right)$
@@
that can be compared with the rate function $i_{t}(q)$ of Eq. 71 concerning the forward propagator alone : the only change is that the typical value where Eq. 111 vanishes is
@@
instead of the typical value $\hat{q}=\lambda^{t}$ of Eq. 47 for the forward propagator.
@@ V Analysis of the backward dynamics for a mixture of two initial conditions $\vec{\sigma}^{[a=1,2]}$
In this section, we consider that the initial condition $P^{ini}(\vec{S})$ is a mixture of two images $\sigma^{[1]}$ and $\sigma^{[2]}$
@@
it is thus useful to introduce the four populations $(\epsilon_{1}=\pm,\epsilon_{2}=\pm)$
@@
The large deviations properties of the joint probability distribution $P^{[Q_{12}]}(Q_{1},Q_{2})$ of Eq. 121 can be analyzed using the Stirling approximation of Eq. 35 with the expression analogous to Eq. 121 to obtain in terms of the three intensive overlaps
@@
where the rate function ${\cal I}^{[q_{12}]}(q_{1},q_{2})$ for the two intensive overlaps $q_{1}$ and $q_{2}$ reads
@@
$\pi_{1}(\vec{S})$ $=\frac{e^{\mu_{T}Nq_{1}}}{e^{\mu_{T}Nq_{1}}+e^{\mu_{T}Nq_{2}}}=\frac{e^{\mu_{T}N(q_{1}-q_{2})}}{e^{\mu_{T}N(q_{1}-q_{2})}+1}\equiv\frac{e^{\mu_{T}Nq_{q}}}{e^{\mu_{T}Nq_{q}}+1}$
@@
will govern the small typical gaussian fluctuations of order $O\left(\frac{1}{\sqrt{N}}\right)$ of $q_{d}$
@@
In conclusion, the ratio of the two weights of Eq. 124 of the two images produced by the backward-generative dynamics
$\frac{\pi_{1}(\vec{S})}{\pi_{2}(\vec{S})}=e^{\mu_{T}Nq_{q}}$ (132)
is determined by the difference $q_{d}=q_{1}-q_{2}$ of the two intensive overlaps, whose large deviations for large $N$ are governed by the rate function $J^{[q_{12}]}(q_{d})$ of Eq. 128 and whose small typical gaussian fluctuations of order $O\left(\frac{1}{\sqrt{N}}\right)$ around the origin $q_{d}=0$ are governed by Eq. 131. As a consequence, the typical scaling of the ratio of Eq. 132 is $e^{\pm\sqrt{N}}$, i.e. the smallest of the two weights will be of order $e^{-\sqrt{N}}$ and thus negligeable with respect to the other complementary weight. Only the initial conditions that have roughly equal intensive overlaps with the two images $q_{d}=q_{1}-q_{2}\simeq 0$ will produce two finite weights, but these initial conditions occur only with a probability of order $\frac{1}{\sqrt{N}}$.
@@ VI Analysis when the initial probability $P^{ini}(\vec{\sigma})$ is given by the Curie-Weiss model
In this section, we focus on the case considered in the previous work [3] where the initial probability $P^{ini}(\vec{\sigma})$ is given by the Curie-Weiss model, whose properties are recalled in the next subsection.
VI.1 Initial condition $P^{ini}(\vec{\sigma})$ given by the Curie-Weiss mean-field ferromagnetic model
@@
In the Curie-Weiss mean-field ferromagnetic model at inverse temperature $\beta$, the Boltzmann probability $P^{ini}(\vec{\sigma})$ of the configuration $\vec{\sigma}$ depends only on its magnetization $M_{0}$
$P^{ini}(\vec{\sigma})=\frac{1}{Z}\sum_{M_{0}}e^{\frac{\beta}{2N}M_{0}^{2}}\delta_{M_{0},\sum_{n}\sigma_{n}}$ (133)
so the probability $p^{ini}(M_{0})$ of the initial magnetization $M_{0}$ itself involves the appropriate binomial coefficient
$p^{ini}(M_{0})=\sum_{\vec{\sigma}}P^{ini}(\vec{\sigma})\delta_{M_{0},\sum_{n}\sigma_{n}}=\frac{1}{Z}e^{\frac{\beta}{2N}M_{0}^{2}}\sum_{\vec{\sigma}}\delta_{M_{0},\sum_{n}\sigma_{n}}=\frac{1}{Z}e^{\frac{\beta}{2N}M_{0}^{2}}\frac{N!}{\left(\frac{N+M_{0}}{2}\right)!\left(\frac{N-M_{0}}{2}\right)!}$ (134)
while the partition function $Z$ ensuring the normalization of probabilities is given by
$Z=\sum_{M_{0}}e^{\frac{\beta}{2N}M_{0}^{2}}\frac{N!}{\left(\frac{N+M_{0}}{2}\right)!\left(\frac{N-M_{0}}{2}\right)!}$ (135)
For large $N$, the probability of the intensive magnetization $m_{0}=\frac{M_{0}}{N}$ displays the large deviation form
$p^{ini}(M_{0}=Nm_{0})\mathop{\propto}_{N\to+\infty}e^{-NI^{ini}(m_{0})}$ (136)
where the rate function $I^{ini}(m_{0})$ is given by the difference
$I^{ini}(m_{0})=V_{\beta}(m_{0})-V^{min}_{\beta}$ (137)
between the effective potential
$V_{\beta}(m_{0})=\frac{(1+m_{0})}{2}\ln(1+m_{0})+\frac{(1-m_{0})}{2}\ln(1-m_{0})-\frac{\beta}{2}m_{0}^{2}$ (138)
and its minimum value $V^{min}_{\beta}$. The two first derivatives of the effective potential $V_{\beta}(m_{0})$
$V_{\beta}^{\prime}(m_{0})$ $=\frac{\ln(1+m_{0})-\ln(1-m_{0})}{2}-\beta m_{0}=\operatorname{arctanh}(m_{0})-\beta m_{0}$
$V_{\beta}^{\prime\prime}(m_{0})$ $=\frac{\frac{1}{1+m_{0}}+\frac{1}{1-m_{0}}}{2}-\beta=\frac{1}{1-m_{0}^{2}}-\beta$ (139)
and the perturbative expansion around the origin $m_{0}=0$
$V_{\beta}(m_{0})=(1-\beta)\frac{m_{0}^{2}}{2}+\frac{m_{0}^{4}}{12}+O(m_{0}^{6})$ (140)
leads to the critical point $\beta_{c}=1$ for the ferromagnetic transition :
(i) in the high temperature phase $\beta<\beta_{c}=1$, the minimum of the potential is at the origin $m_{0}=0$.
(ii) in the low temperature phase $\beta<\beta_{c}=1$, the origin $m_{0}=0$ is instead a local maximum, while the solution of $0=V_{\beta}^{\prime}(m_{0})=\operatorname{arctanh}(m_{0})-\beta m_{0}$ leads to the two symmetric minima at [±$[\pm m_{F}(\beta)\neq 0$ corresponding to the intensive spontaneous magnetization of the ferromagnetic phase, where the rate function of Eq. 137 vanishes
$I^{ini}(m_{0})=0\text{ for }\ m_{0}=\pm m_{F}(\beta)\ \ \ \text{with }\ m_{F}(\beta)=\tanh(\beta m_{F}(\beta))$ (141)
@@ VI.2 Properties of the forward dynamics for the magnetization $M$
The forward dynamics in the space of configurations $\vec{S}$ can be projected onto the forward dynamics for the magnetization $M$
@@
where the forward propagator $p_{t}(M|M_{0})$ corresponds to the sum over the overlap $Q$ of the joint distribution $P^{[M_{0}]}_{t}(M,Q)$ analyzed in Eqs 25 and 32
@@
The large deviations of the forward solution of Eq. 142
[…]
are then governed by the rate function $I_{t}(m)$ corresponding to the following optimization over the initial magnetization $m_{0}$
$I_{t}(m)=\min_{m_{0}}\left[I_{t}^{[m_{0}]}(m)+I^{ini}(m_{0})\right]$ (146)
@@ VI.3 Properties of the backward dynamics for the magnetization $M$
Since the uniform steady stae $P_{*}(\vec{S}(T))=\frac{1}{2^{N}}$ in the configuration space translates into the binomial distribution $p_{*}(M_{T})$ for the magnetization $M_{T}$
@@
$b^{Gene}_{t}(M)$ $=\sum_{M_{T}}p_{*}(M_{T})\frac{p_{T-t}(M_{T}|M)p_{t}(M|M_{0})p^{ini}(M_{0})}{\sum_{M_{0}^{\prime}}p_{T}(M_{T}|M_{0}^{\prime})p^{ini}(M_{0}^{\prime})}$ (148)
$=p_{t}(M|M_{0})p^{ini}(M_{0})\sum_{M_{T}}2^{-N}\frac{N!}{\left(\frac{N+M_{T}}{2}\right)!\left(\frac{N-M_{T}}{2}\right)!}\times\frac{p_{T-t}(M_{T}|M)}{\sum_{M_{0}^{\prime}}p_{T}(M_{T}|M_{0}^{\prime})p^{ini}(M_{0}^{\prime})}$
@@
Plugging the large deviations of the binomial coefficient
$2^{-N}\frac{N!}{\left(\frac{N+M_{T}}{2}\right)!\left(\frac{N-M_{T}}{2}\right)!}\mathop{\propto}_{N\to+\infty}e^{\displaystyle-N\left[\ln 2+\sum_{\eta=\pm 1}\frac{1+\eta m_{T}}{2}\ln\left(\frac{1+\eta m_{T}}{2}\right)\right]}$ (150)
and of the other factors discussed previously into Eq. 148 yields the large deviations of the intensive magnetization $m$ generated by the backward dynamics at time $t$
$b^{Gene}_{t}(M=Nm)\mathop{\propto}_{N\to+\infty}e^{\displaystyle-N\left(I_{t}^{[m_{0}]}(m)+I^{ini}(m_{0})\right)}\int m_{T}e^{\displaystyle-N\left[\ln 2+\sum_{\eta=\pm 1}\frac{1+\eta m_{T}}{2}\ln\left(\frac{1+\eta m_{T}}{2}\right)+I_{T-t}^{[m]}(m_{T})-I_{T}(m_{T})\right]}$
$\mathop{\propto}_{N\to+\infty}e^{\displaystyle-N\left(I_{t}^{[m_{0}]}(m)+I^{ini}(m_{0})\right)+\min_{m_{T}}\left[\ln 2+\sum_{\eta=\pm 1}\frac{1+\eta m_{T}}{2}\ln\left(\frac{1+\eta m_{T}}{2}\right)+I_{T-t}^{[m]}(m_{T})-I_{T}(m_{T})\right]}$ (151)
So one needs to optimize over $m_{T}$ the function in the square parenthesis that involves the rate function $I_{T-t}^{[m]}(m_{T})$ discussed around Eqs 76 88 as well as the rate function $I_{T}(m_{T})$ of Eq. 146 corresponding to an optimization over $m_{0}$
$I_{T}(m)=\min_{m_{0}}\left[I_{T}^{[m_{0}]}(m_{T})+I^{ini}(m_{0})\right]$ (152)
In summary, when the initial probability $P^{ini}(\vec{\sigma})$ is given by the Curie-Weiss model, one can analyze the large deviations properties of the intensive magnetization both for the forward dynamics and the backwards dynamics, even if the explicit implementation of the remaining optimizations over the initial magnetization $m_{0}$ and over the final magnetization $m_{T}$ as a function of the inverse temperature $\beta$ of the initial Curie-Weiss distribution go beyond the goals of the present work.

### v2

@@ IV.3 Large deviations of the backward-generated distribution of the intensive overlap $q$ for large $N$
For large $N$, using the Stirling approximation of Eq. 28, one obtains that the large deviations of the intensive overlap $q=\frac{Q}{N}$ in the backward-generated distribution of Eq. 96
@@
that has the same form as the rate function $i_{t}(q)$ of Eq. 30 concerning the forward propagator alone: the only change is that the forward typical value $\hat{q}=\lambda^{t}$ has been replaced by the backward typical value
@@
IV.4 Discussion
In summary, when the initial condition reduces to a single image, the backward generative dynamics has very simple properties for any time $t$ in terms of the extensive overlap $Q$ for any finite number $N$ of pixels, as well as in terms of the intensive overlap $q=\frac{Q}{N}$ when one focuses on large deviations for large $N\to+\infty$.
@@ V Analysis of the backward dynamics for a mixture of two initial conditions $\vec{\sigma}^{[a=1,2]}$
In this section, we consider the case where the initial condition $P^{ini}(\vec{S})$ is a mixture of two images $\sigma^{[1]}$ and $\sigma^{[2]}$
@@
it is useful to introduce the four populations $(\epsilon_{1}=\pm,\epsilon_{2}=\pm)$
@@
The large deviations properties of the joint probability distribution $P^{[Q_{12}]}(Q_{1},Q_{2})$ of Eq. 108 in terms of the three intensive overlaps
@@
can be analyzed as in the analog computation described in the Appendix section A.1 to obtain the rate function
@@
$\pi_{1}(\vec{S})$ $=\frac{e^{\mu_{T}Nq_{1}}}{e^{\mu_{T}Nq_{1}}+e^{\mu_{T}Nq_{2}}}=\frac{e^{\mu_{T}N(q_{1}-q_{2})}}{e^{\mu_{T}N(q_{1}-q_{2})}+1}\equiv\frac{e^{\mu_{T}Nq_{d}}}{e^{\mu_{T}Nq_{d}}+1}$
@@
will govern the Gaussian typical fluctuations of order $O\left(\frac{1}{\sqrt{N}}\right)$ of $q_{d}$
@@
In conclusion, the ratio of the two weights of Eq. 111 for the two images produced by the backward-generative dynamics
$\frac{\pi_{1}(\vec{S})}{\pi_{2}(\vec{S})}=e^{\mu_{T}Nq_{d}}$ (119)
is determined by the difference $q_{d}=q_{1}-q_{2}$ between the two intensive overlaps, whose large deviations for large $N$ are governed by the rate function $J^{[q_{12}]}(q_{d})$ of Eq. 115 and whose Gaussian typical fluctuations of order $O\left(\frac{1}{\sqrt{N}}\right)$ around the origin $q_{d}=0$ are governed by Eq. 118. As a consequence, the typical scaling of the ratio of Eq. 119 is $e^{\pm\sqrt{N}}$, i.e. the smallest of the two weights will be of order $e^{-\sqrt{N}}$ and thus negligible with respect to the other complementary weight. Only the initial conditions that have roughly equal intensive overlaps with the two images $q_{d}=q_{1}-q_{2}\simeq 0$ will produce two finite weights, but these initial conditions occur only with a probability of order $\frac{1}{\sqrt{N}}$.
V.4 Discussion
In summary, when the initial condition reduces to a mixture of two images with equal weights, the backward generative dynamics starting at configuration $\vec{S}(T)$ will reproduce the two images with unequal weights $\pi_{1,2}(\vec{S}(T))$ that are determined by the overlaps, and whose statistical properties can be analyzed explicitly both for any finite number $N$ of pixels and in the limit of large $N\to+\infty$.
@@ VI Analysis when the initial probability $P^{ini}(\vec{\sigma})$ is given by the Curie-Weiss model
In this section, we focus on the case where the initial probability $P^{ini}(\vec{\sigma})$ is given by the Curie-Weiss mean-field ferromagnetic model: while the canonical ensemble at inverse temperature $\beta$ has been considered in the previous work [3], here we will focus instead on the microcanonical ensemble in order to have an equivalent of the manifold-hypothesis as discussed in the next subsection.
VI.1 Equivalent of the manifold-hypothesis for discrete models involving $N$ classical spins $S_{n}=\pm 1$
In the field of generative diffusion models, the manifold-hypothesis (see [96, 97, 12, 13] and references therein) means that the empirical distribution constructed from the set of interesting images actually lives on a low-dimensional manifold within the global manifold containing all possible images, and the real goal is in fact to learn this low-dimensional manifold. Indeed in practical applications, one wishes to avoid the perfect memorization of the initial data and to generate instead new images that are close but nevertheless different from the initial data (see [5, 7, 13, 98, 99] and references therein).
When translated for our present discrete models involving $N$ classical spins $S_{n}=\pm 1$ where there are $2^{N}$ possible images, the analog of the low-dimensional manifold is a subset of images whose size grows less rapidly than $2^{N}$, so that the meaningful images of interest become ’rare’ for large $N$. Let us give two simple examples.
VI.1.1 Uniform distribution $P^{ini}_{M_{F}}(\vec{\sigma})$ among the configurations $\vec{\sigma}$ with a given magnetization $M_{F}$
The simplest example is when the configuration $\vec{\sigma}$ is drawn uniformly within the set of configurations with a given fixed magnetization $M_{F}$
$P^{ini}_{M_{F}}(\vec{\sigma})=\frac{1}{\Omega_{N}(M_{F})}\delta_{M_{F},\sum_{n}\sigma_{n}}$ (120)
whose size $\Omega_{N}(M_{F})$ is given by the binomial coefficient
$\Omega_{N}(M_{F})\equiv\sum_{\vec{\sigma}}\delta_{M_{F},\sum_{n}\sigma_{n}}=\frac{N!}{\left(\frac{N+M_{F}}{2}\right)!\left(\frac{N-M_{F}}{2}\right)!}$ (121)
so that the ratio with the total number $2^{N}$ of possibles images
$\frac{\Omega_{N}(M_{F}=Nm_{F})}{2^{N}}\mathop{\simeq}_{N\to+\infty}e^{-N\left[\frac{(1+m_{F})}{2}\ln(1+m_{F})+\frac{(1-m_{F})}{2}\ln(1-m_{F})\right]}$ (122)
is exponentially small in $N$ as long as the intensive magnetization $m_{F}=\frac{M_{F}}{N}\in[-1,+1]$ does not vanish $m_{F}\neq 0$.
VI.1.2 Curie-Weiss mean-field ferromagnetic model in the microcanonical ensemble where the energy $E_{F}=-\frac{M_{F}^{2}}{2N}$ is fixed
@@ VI.1 Equivalent of the manifold-hypothesis for discrete models involving $N$ classical spins $S_{n}=\pm 1$
In the Curie-Weiss mean-field ferromagnetic model, the energy $E(\vec{\sigma})$ of the configuration $\vec{\sigma}$ depends only on its magnetization
$E(\vec{\sigma})=-\frac{1}{2N}\left(\sum_{n=1}^{N}\sigma_{n}\right)^{2}$ (123)
So in the microcanonical ensemble where the configuration $\vec{\sigma}$ is drawn uniformly among the configurations with a given non-vanishing energy $E_{F}=-\frac{M_{F}^{2}}{2N}$ that is associated to the two opposite magnetizations (±$(\pm M_{F})$, the distribution reduces to the mixture of the two fixed-magnetization distributions $P^{ini}_{\pm M_{F}}(\vec{\sigma})$ of Eq. 120
$P^{ini}_{E_{F}=-\frac{M_{F}^{2}}{2N}}(\vec{\sigma})=\frac{1}{2}\left(P^{ini}_{M_{F}}(\vec{\sigma})+P^{ini}_{-M_{F}}(\vec{\sigma})\right)=\frac{1}{2\Omega_{N}(M_{F})}\left(\delta_{M_{F},\sum_{n}\sigma_{n}}+\delta_{-M_{F},\sum_{n}\sigma_{n}}\right)$ (124)
@@ VI.2 Properties of the forward dynamics for the magnetization $M$
The forward dynamics in the space of configurations $\vec{S}$ can be projected onto the forward dynamics for the magnetization $M$ alone
@@
where the initial distribution $p^{ini}(M_{0})$ reduces to a single delta function for the case of Eq. 120
$p^{ini}_{M_{F}}(M_{0})=\delta_{M_{0},M_{F}}$ (126)
or to the mixture of two delta functions for the case of Eq. 124
$p^{ini}_{E_{F}=-\frac{M_{F}^{2}}{2N}}(M_{0})=\frac{1}{2}\left(\delta_{M_{0},M_{F}}+\delta_{M_{0},-M_{F}}\right)$ (127)
The forward propagator $p_{t}(M|M_{0})$ governing Eq. 125 corresponds to the sum over the overlap $Q$ of the joint distribution $P^{[M_{0}]}_{t}(M,Q)$ analyzed in Eqs 32 and 39
@@ VI.3 Properties of the backward dynamics for the magnetization $M$
Since the uniform steady state $P_{*}(\vec{S}(T))=\frac{1}{2^{N}}$ in the spin configuration space translates into the binomial distribution $p_{*}(M_{T})$ for the global magnetization $M_{T}$
@@
$b^{Gene}_{t}(M)$ $=p_{t}(M)\sum_{M_{T}}p_{*}(M_{T})\frac{p_{T-t}(M_{T}|M)}{p_{T}(M_{T})}$ (131)
$=p_{t}(M)\sum_{M_{T}}2^{-N}\frac{N!}{\left(\frac{N+M_{T}}{2}\right)!\left(\frac{N-M_{T}}{2}\right)!}\times\frac{p_{T-t}(M_{T}|M)}{p_{T}(M_{T})}$
@@
and with the large deviations of the steady distribution $p_{*}(M_{T})$ of Eq. 130
$p_{*}(M_{T})=2^{-N}\frac{N!}{\left(\frac{N+M_{T}}{2}\right)!\left(\frac{N-M_{T}}{2}\right)!}\mathop{\propto}_{N\to+\infty}e^{-N\left[\frac{1+m_{T}}{2}\ln\left(1+m_{T}\right)+\frac{1-m_{T}}{2}\ln\left(1-m_{T}\right)\right]}$ (133)
VI.3.1 Application to the fixed-magnetization ensemble $p^{ini}_{M_{F}}(M_{0})=\delta_{M_{0},M_{F}}$
For the fixed-magnetization case of Eq. 126, the forward solution $p_{t}(M)$ of Eq. 125 reduces to the forward propagator $p_{t}(M|M_{F})$ whose large deviations properties of Eq. 129 are governed by the rate function $I_{t}^{[m_{F}]}(m)$
$p_{t}(M)=p_{t}(M|M_{F})\mathop{\propto}_{N\to+\infty}e^{-NI_{t}^{[m_{F}]}(m)}$ (134)
The backward-generated distribution $b^{Gene}_{t}(M)$ of Eq. 131 becomes
$b^{Gene}_{t}(M)$ $=p_{t}(M|M_{F})\sum_{M_{T}}p_{*}(M_{T})\frac{p_{T-t}(M_{T}|M)}{p_{T}(M_{T}|M_{F})}$ (135)
$=p_{t}(M|M_{F})\sum_{M_{T}}2^{-N}\frac{N!}{\left(\frac{N+M_{T}}{2}\right)!\left(\frac{N-M_{T}}{2}\right)!}\times\frac{p_{T-t}(M_{T}|M)}{p_{T}(M_{T}|M_{F})}$
with the large deviations behavior
[…]
$\mathop{\simeq}_{N\to+\infty}e^{\displaystyle-N\left(I_{t}^{[m_{F}]}(m)+J_{t}^{[m_{F}]}(m)\right)}$
where the supplementary contribution $J_{t}^{[m_{F}]}(m)$ with respect to $I_{t}^{[m_{F}]}(m)$ governing the forward solution of Eq. 134
$J_{t}^{[m_{F}]}(m)\equiv\min_{m_{T}\in[-1,1]}\left[\frac{1+m_{T}}{2}\ln\left(1+m_{T}\right)+\frac{1-m_{T}}{2}\ln\left(1-m_{T}\right)-I_{T-t}^{[m]}(m_{T})+I_{T}^{[m_{F}]}(m_{T})\right]$ (137)
requires the optimization over $m_{T}\in[-1,1]$ of the function in the square parenthesis that involves the difference between the two rate functions $I_{T-t}^{[m]}(m_{T})$ and $I_{T}^{[m_{F}]}(m_{T})$.
VI.3.2 Application to the Curie-Weiss microcanonical ensemble $p^{ini}_{E_{F}=-\frac{M_{F}^{2}}{2N}}(M_{0})=\frac{1}{2}\left(\delta_{M_{0},M_{F}}+\delta_{M_{0},-M_{F}}\right)$
For the Curie-Weiss microcanonical ensemble of Eq. 127, the forward solution $p_{t}(M)$ of Eq. 125 reduces to the mixture of two forward propagators whose large deviations properties of Eq. 129 are governed by the two rate functions $I_{t}^{[\pm m_{F}]}(m)$
$p_{t}(M)$ $=\frac{p_{t}(M|M_{F})+p_{t}(M|-M_{F})}{2}\mathop{\propto}_{N\to+\infty}\frac{e^{-NI_{t}^{[m_{F}]}(m)}+e^{-NI_{t}^{[-m_{F}]}(m)}}{2}$ (138)
$\mathop{\propto}_{N\to+\infty}e^{-N\min\left[I_{t}^{[m_{F}]}(m);I_{t}^{[-m_{F}]}(m)\right]}\mathop{\propto}_{N\to+\infty}e^{-NI_{t}^{[m_{F}{\rm sgn}(m)]}(m)}$
where the minimum between the two rate functions $I_{t}^{[\pm m_{F}]}(m)$ is obtained when the two magnetizations have the same sign.
The backward-generated distribution $b^{Gene}_{t}(M)$ of Eq. 131 becomes
$b^{Gene}_{t}(M)$ $=\bigg{(}p_{t}(M|M_{F})+p_{t}(M|-M_{F})\bigg{)}\sum_{M_{T}}p_{*}(M_{T})\frac{p_{T-t}(M_{T}|M)}{p_{T}(M_{T}|M_{F})+p_{T}(M_{T}|-M_{F})}$ (139)
$=\bigg{(}p_{t}(M|M_{F})+p_{t}(M|-M_{F})\bigg{)}\sum_{M_{T}}2^{-N}\frac{N!}{\left(\frac{N+M_{T}}{2}\right)!\left(\frac{N-M_{T}}{2}\right)!}\times\frac{p_{T-t}(M_{T}|M)}{p_{T}(M_{T}|M_{F})+p_{T}(M_{T}|-M_{F})}$
with the large deviations behavior for the intensive magnetization $m$
[…]
$\mathop{\simeq}_{N\to+\infty}e^{\displaystyle-N\left(I_{t}^{[m_{F}{\rm sgn}(m)]}(m)+K_{t}^{[m_{F}]}(m)\right)}$
where the supplementary contribution $K_{t}^{[m_{F}]}(m)$ with respect to $I_{t}^{[m_{F}{\rm sgn}(m)]}(m)$ governing the forward solution of Eq. 138
$K_{t}^{[m_{F}]}(m)\equiv\min_{m_{T}\in[-1,1]}\left[\frac{1+m_{T}}{2}\ln\left(1+m_{T}\right)+\frac{1-m_{T}}{2}\ln\left(1-m_{T}\right)-I_{T-t}^{[m]}(m_{T})+I_{T}^{[m_{F}{\rm sgn}(m_{T})]}(m_{T})\right]$ (141)
requires the optimization over $m_{T}\in[-1,1]$ of the function in the square parenthesis that involves the difference between the two rate functions $I_{T-t}^{[m]}(m_{T})$ and $I_{T}^{[m_{F}{\rm sgn}(m_{T})]}(m_{T})$.